OpenStudy (anonymous):

How do I calculate the derivative of 2x^3 + 12x^2 - 30x +10? thanks.

OpenStudy (anonymous):

power rule: 6x^2+24x-30

OpenStudy (anonymous):

differentiate term by term using the general form if f\9) = ax^n then f'(x) = anx^(n-1)

OpenStudy (anonymous):

ax^n =a*nx^n-1

OpenStudy (anonymous):

if f(x) i meant to write!!

OpenStudy (anonymous):

Great, quick answers. Thanks. Could I use the definition and determine it by doing [2(x+h)^3+ 12(x+h)^2 ... etc.? And the power series method is just easier? Or is the definition not appropriate in this context?

OpenStudy (anonymous):

fine the Ybbutton witha x above it

OpenStudy (anonymous):

the Y above* it

OpenStudy (anonymous):

you could use the definition but the power method is valid here

OpenStudy (anonymous):

Perfect. Thanks.